Two standard dice are rolled. What is the probability of the sum on the two dice being less than 10.
Answer by stanbon(75887)

Two standard dice are rolled. What is the probability of the sum on the two dice being less than 10.
---------------
Draw the sample space which is
a 6 by 6 square with the sums listed:
2,3,4,5,6,7
3,4,5,6,7,8
4,5,6,7,8,9
...
...
7,8,9,10,11,12,
----------------------
Count the sums that are less than 10: 30
----------------------
P(sum less than 10) = 30/36 = 5/6
==========================================
Cheers,
Stan H.